OpenStudy (anonymous):

Solve the equation.. 2(x-3)= 5 divided by 2

OpenStudy (anonymous):

2(x-3)= 5 /2 Distribute. 2x - 6 = 5 /2 Combine like terms. 2x - 6 = 5 /2 2x - 6 = 2.5 2x = 2.5 + 6 2x = 8.5 x = 4.25 or as a fraction: 4 1/4

OpenStudy (anonymous):

okay thanks much @jazy

OpenStudy (anonymous):

No problem. (: